OpenStudy (anonymous):

use suitable substitution and table of integral to evaluate ::: integral [ arctan (x/2) ]/ [x^2 +4] dx ans: 1/4 ( arctan (x/2)^2 ) +c

OpenStudy (thomas5267):

Is this the integral? \[ \int\frac{\arctan\left(\frac{x}{2}\right)}{x^2+4}\,dx \]

OpenStudy (thomas5267):

Hint: Integration by parts.